# Security Code Review by Large Language Models

### v2

@@ 7. Conclusions
In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems present in responses generated by this LLM and performed regression analysis to explore factors influencing its performance. Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to the state-of-the-art static analysis tools. However, their application remains limited. (3) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code that has fewer tokens, containing functional logic and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.

### v3

@@ 7. Conclusions and Future Work
In our study, we conducted a comprehensive analysis of the practicality of current popular LLMs in the context of security code review. We investigated the capabilities of 6 pre-trained LLMs in detecting security defects similar to those identified by human reviewers during real-world code reviews. For the best-performing LLM, we conducted a linguistic analysis of the quality problems in the responses generated by this LLM. We performed a regression analysis to explore factors influencing its performance.
Our main findings are: (1) In security code review tasks, LLMs demonstrate better capabilities compared to state-of-the-art static analysis tools. However, their application remains limited. (2) GPT-4 performs best among all LLMs when provided with a CWE list for reference. (3) During code review, GPT-4 is more effective in identifying security defects in code files with fewer tokens, containing functional logic, and written by developers with less involvement in the project. (4) The most prevalent quality problems in the responses generated by GPT-4 are verbosity and non-compliance with task requirements.
We plan to extend this work in the next step by incorporating code reviews from proprietary software projects into our dataset and attempting enhancement strategies such as few-shot learning and fine-tuning. In addition to applying LLMs to help reviewers detect security defects during code review, we also plan to analyze the ability of LLMs to fix security defects in code so that developers can comprehensively evaluate the potential of LLMs in the task of secure code review.
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.